write an equation of the line that passes through (2,3) & (-3,4)
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

write an equation of the line that passes through (2,3) & (-3,4)
-----
slope = m = (4-3)/(-3-2) = -1/5
-------------------------------
Form: y = mx + b
Substitute to solve for "b":
3 = (-1/5)2 + b
3 = -2/5 + b
b = 15/5 + 2/5
b = 17/5
=========
Cheers,
Stan H.

Answer by mananth(16949)

(2,3) & (-3,4)
slope m = y2-y1/x2-x1
...
m= 4-3 / -3-2
m=-1/5
..
y = mx+b
point(2,3)
find b
3=-1/5*2+b
3+2/5 =b
17/5 =b
...
y= mx + b
y=(-1/5)x+17/5
y=-x/5+17/5
...
m.ananth@hotmail.ca